# Distributions of Ulam Words up to Length 30

### v1

@@ Abstract.
We further explore the notion of Ulam words considered by Bade, Cui, Labelle, and Li. We find that when interpreted as integers in a natural way, Ulam words appear to follow a new, unexplained distribution. Gaps between words and words of special type also reveal remarkable structure. By substantially increasing the number of computed terms, we are also able to sharpen some of the conjectures made by Bade et al.
@@ 1. Introduction:
In their 2020 paper[1], Bade, Cui, Labelle, and Li introduced the notion of Ulam words, defined as follows. Consider the free semigroup $S[\{0,1\}]$ on two generators 00 and $1$. We say that 00 and $1$ are Ulam and then define all other Ulam words inductively: a word $w\neq 0,1$ is Ulam if and only if there exist exactly one pair of Ulam words $u_{1}\neq u_{2}$ such that $w=u_{1}{}^{\frown}u_{2}$. (Here, denotes concatenation.) We shall denote the entire set of Ulam words as $\mathscr{U}$, and Ulam words of length $n$ by $\mathscr{U}_{n}$. It is easy to check that:
@@
Bade et al. computed all Ulam words up to length $24$—using an algorithm described in Section 6, we were able to compute up to length $30$. While this might appear as a small improvement at first glance, because the number of Ulam words of length $n$ (almost) doubles on each iteration, in reality, this represents nearly $60$ times as much data. One consequence of this additional data is that we can better project the growth rate of the density
$\rho(n):=\frac{\#\mathscr{U}_{n}}{2^{n}}$
as a function of $n$—see Section 4 for details, and Figure 1 for an illustration. It has also provided us with substantial numerical evidence for a new class of conjectures regarding the distribution of Ulam words, which we shall describe below.
@@
We will discuss this construction more precisely in Section 3, but, briefly, the discrete Sierpiński triangle is an approximation to the standard Sierpiński triangle. It can be constructed either iteratively (as in Figure 2) or by coloring Pascal’s triangle by parity.
@@
Figure 2. Visual of the first 4 steps of constructing discrete Sierpiński triangle.
On the other hand, we also have a novel way of considering Ulam words, by interpreting them as integers. Observe that there exists a natural map $\pi:S(\{0,1\})\rightarrow\mathbb{Z}_{\geq 0}$ via interpreting a word as the binary representation of an integer. In general, this map is not injective—for example, $\pi(0)=\pi(00)=\pi(000)=0$. However, if we restrict it to words of a fixed length, then it is. In particular, the restrictions $\pi:\mathscr{U}_{n}\rightarrow\mathbb{Z}\cap[0,2^{n-1}]$ are injective maps. This gives a natural ordering on $\mathscr{U}_{n}$ and allows us to ask questions about equidistribution. For instance: do Ulam words equidistribute modulo $N$?
@@
As we discuss in Section 5, while this conjecture is consistent with the available data, it is somewhat surprising. For one thing, $\rho_{5,6}(1)=\rho_{5,6}(2)=\rho_{5,6}(3)=0$, and it takes some time before it appears to start to converge to $1/6$.
This conjecture can be viewed as saying that Ulam words equidistribute with respect to the non-archimedean valuations (essentially, modulo arbitrarily large prime powers). We might also ask whether they equidistribute with respect to the archimedean valuation, which we choose to measure as follows. Consider normalizing each $\mathscr{U}_{n}$ such that it is within the interval $[0,1]$—i.e. we consider instead $2^{-n}\mathscr{U}_{n}$. Then we can apply the usual notion of the equidistribution of a sequence, which is that in the limit the proportion of points in an interval is the length of the interval. That is, is it true that for all $0\leq c<d\leq 1$,
$\lim_{n\rightarrow\infty}\frac{\#\left\{w\in\mathscr{U}_{n}\middle|2^{-n}w\in[c,d]\right\}}{\#\mathscr{U}_{n}}=d-c?$
It would appear not, as Figure 3 illustrates! It is clear that if $c,d\approx 1/2$ then there will be significantly fewer points in the interval than expected. Nevertheless, it appears that the limit does approach some distribution, even if cannot determine what it is precisely. To be concrete, we tentatively make the following conjecture.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 3. From left to right, top to bottom: histograms of words of length $25,26,27,28,29,30$.
Conjecture 1.2.
Given a positive integer $n$, define a probability measure $\delta_{\mathscr{U}_{n}}:[0,1]\rightarrow[0,\infty)$ by
$\delta_{\mathscr{U}_{n}}(S)=\frac{\#\left(2^{-n}\mathscr{U}_{n}\cap S\right)}{\#\mathscr{U}_{n}}.$
As $n\rightarrow\infty$, $\delta_{\mathscr{U}_{n}}$ converges weakly to some probability distribution $\delta_{\mathscr{U}}:[0,1]\rightarrow[0,\infty)$.
There is a related conjecture that we can make with greater confidence, which is that if we look at the distribution of the gaps between consecutive Ulam words, this converges as $n\rightarrow\infty$.
Conjecture 1.3.
$p_{n}:\mathbb{Z}_{\geq 1}$ $\rightarrow[0,\infty)$
$g$$\mapsto\frac{\#\left\{i\middle|u_{i+1}-u_{i}=g\right\}}{k_{n}-1}.$
This has a natural interpretation as a probability measure. As $n\rightarrow\infty$, the functions $p_{n}$ converge pointwise to a probability measure $p:\mathbb{Z}_{\geq 1}\rightarrow[0,\infty)$.
Illustrations of the probability measures $p_{n}$ and some discussion about their properties can be found in Section 4.
@@ 2. Definitions and Visualizations
We start with some basic definitions and constructions. Given a word $w\in S[{0,1}]$, we define its complement $\hat{w}$ to be the word with every instance of 00 replaced with a $1$, and vice versa. We also define the reverse $\overline{w}$, which is the word obtained by reversing the order of the letters. It was shown by Bade et al. [1] that $w\in\mathscr{U}$ if and only if $\hat{w}\in\mathscr{U}$, if and only if $\overline{w}\in\mathscr{U}$.
In order to better visualize the set $\mathscr{U}$, we made use of heat maps, which depict each Ulam word as a colored bar and stacks all of the words vertically—that is, for a given word, a 00 corresponds to a rectangle of one color, and a $1$ corresponds to a rectangle of a second color. An example is provided in Figure 4. In general, we abridge such diagrams: we created figures only using all the Ulam words that started with zero, since Ulam words are closed under complements.
Figure 4. Example of how an Ulam word appears in a heatmap. Heat map shows all Ulam words of length 6 without forced ordering.
Patterns in Ulam words become apparent in the heat maps if we impose the ordering discussed in the introduction: we say that $w\leq w^{\prime}$ if and only if $\pi(w)\leq\pi(w^{\prime})$. See Figure 5. Using this way of visualizing Ulam words allows us to easily see that there is both a clear binary tree structure that governs the existence of Ulam words, as well as a chaotic element to the set where the binary tree breaks down.
@@
We can be more specific about our meaning regarding this breakdown: since Ulam words are preserved under the reverse map, this is equivalent to saying that for any $n$ there exists $n>\ell_{n}>0$ such that all possible subwords of length $\ell_{n}$ occur as the final $\ell_{n}$ characters of words in $\mathscr{U}_{n}$. In turn, that is equivalent to saying that the quotient map $\pi(\mathscr{U}_{n})\rightarrow\mathbb{Z}/2^{\ell_{n}}\mathbb{Z}$ is surjective. Our observation is that $\ell_{n}$ appears to increase as a function of $n$, albeit not very quickly—see Figure 6.
@@
Assuming that Conjecture 1.1 is true, then it would follow immediately that $\ell_{n}\rightarrow\infty$ simply by considering the case where $N=2^{\ell_{n}}$—indeed, the heat maps were the original impetus for our equidistribution conjectures. On the other hand, the “chaotic” latter half of the heat map is more of a mystery.
3. Unconditional Results
Bade et al. proved various results regarding Ulam words containing certain patterns [1]. We offer a similar collection of results. We begin with a couple of intermediate results that to allow us to prove Theorem 1.1. The first two are due to Bade et al.
Theorem 3.1 (Theorem 3.4 in [1]).
A word of the form $0^{a}10^{b}$ is Ulam if and only if
$\left(\begin{smallmatrix}a+b\\ a\end{smallmatrix}\right)=1\mod 2.$
Theorem 3.2 (Theorem 3.5 in [1]).
A word of the form $0^{a}1^{2}0^{b}$ is Ulam if and only if the length of the word is odd (that is, $a+b=1\pmod{2}$).
@@ 3. Unconditional Results
  1. (1)
    $1{}^{\frown}1^{2}0^{n-3}$ and
  2. (2)
    $1^{3}0^{n-4}{}^{\frown}0$,
Lemma 3.4.
Let $u\in S[\{0,1\}]$ be a word of length $n\geq 5$ with exactly four 1’s such that the 1’s are consecutive. The word $u\in\mathscr{U}$ if and only if $n=1\pmod{4}$.
Proof.
We will use proof by induction, where the base cases $n=5,6,7,8$ can be verified directly. Assume the statement holds for all words of length strictly less than $n$, and consider the word $u=0^{k}1^{4}0^{l}$ of length $n$, where since $n\geq 9$ at least one of $k\geq 3$ or $l\geq 3$. By applying the reverse map if necessary, we may assume that $k\geq 3$.

### v2

@@ Abstract.
We further explore the notion of Ulam words considered by Bade, Cui, Labelle, and Li, giving some lower bounds on how many there are of a given length. Gaps between words and words of special type also reveal remarkable structure. By substantially increasing the number of computed terms, we are also able to sharpen some of the conjectures made by Bade et al.
@@ 1. Introduction:
In their 2020 paper[1], Bade, Cui, Labelle, and Li introduced the notion of Ulam words, defined as follows. Consider the free semigroup $S[\{0,1\}]$ on two generators 00 and $1$. We say that 00 and $1$ are Ulam and then define all other Ulam words inductively: a word $w\neq 0,1$ is Ulam if and only if there exists exactly one pair of Ulam words $u_{1}\neq u_{2}$ such that $w=u_{1}{}^{\frown}u_{2}$. (Here, denotes concatenation.) We shall denote the entire set of Ulam words as $\mathscr{U}$, and Ulam words of length $n$ by $\mathscr{U}_{n}$. It is easy to check that:
@@
Bade et al. computed all Ulam words up to length $24$; we were able to compute up to length $30$. While this might appear as a small improvement at first glance, because the number of Ulam words of length $n$ appears to (almost) double on each iteration, in reality, this represents nearly $60$ times as much data.
Of course, it is an open question whether #$\#\mathscr{U}_{n}$ grows exponentially. The best general bound that we can prove is linear, mainly using explicit constructions of words from Bade et al.
Theorem 1.1.
For all $n\geq 6$, #$\#\mathscr{U}_{n}\geq 2n+4$.
However, we are able to demonstrate that there is a subsequence of Ulam words that grows exponentially, using a completely different argument.
Theorem 1.2.
There exists $1<\alpha_{0}\leq 2$ such that for all $1<\alpha<\alpha_{0}$,
$\limsup_{n\rightarrow\infty}\frac{\#\mathscr{U}_{n}}{\alpha^{n}}=\infty.$
Concretely,
$\alpha_{0}=\left(\frac{101847671}{31}\right)^{1/5}\approx 1.648996$
suffices.
We give proofs of both of these theorems in Section 3. Unfortunately, both of these results are still quite far from what is conjectured to hold. To wit, define the density
$\rho(n):=\frac{\#\mathscr{U}_{n}}{2^{n}};$
Bade et al. conjectured that $\rho(n)\rightarrow r$ for some $10<r<$ 1 (see Conjecture 3.10 in [1]); with our enlarged data set, we instead posit something a little stranger.
Conjecture 1.1.
$\rho(n)=\Theta(n^{-3/10})$.
This conjecture is supported by the numerical evidence—see Figures 1 and 2 for an example—but it also has ties to another conjecture involving the average gap between Ulam words, which we shall describe below. In any case, observe that if either conjecture is correct, the number of Ulam words grows only very slightly slower than $2^{n}$.
@@
$\begin{array}[]{l|l}n&\mathscr{U}_{n}\\ \hline\cr 13&1916\\ 14&3812\\ 15&7772\\ 16&14822\\ 17&29368\\ 18&58478\end{array}$
$\begin{array}[]{l|l}n&\mathscr{U}_{n}\\ \hline\cr 19&114300\\ 20&225166\\ 21&441724\\ 22&876238\\ 23&1717748\\ 24&3406884\end{array}$
$\begin{array}[]{l|l}n&\mathscr{U}_{n}\\ \hline\cr 25&6720784\\ 26&13303332\\ 27&26273948\\ 28&52010642\\ 29&102933200\\ 30&203695342\end{array}$
Figure 2. The exact counts for #$\#\mathscr{U}_{n}$ for $13\leq n\leq 30$.
@@
We will discuss this construction more precisely in Section 4, but briefly, the discrete Sierpiński triangle is an approximation to the standard Sierpiński triangle. It can be constructed either iteratively (as in Figure 3) or by coloring Pascal’s triangle by parity.
@@
Figure 3. Visual of the first 4 steps of constructing the discrete Sierpiński triangle.
On the other hand, we also have a novel way of considering Ulam words by interpreting them as integers. Observe that there exists a natural map $\pi:S[\{0,1\}]\rightarrow\mathbb{Z}_{\geq 0}$ via interpreting a word as the binary representation of an integer. In general, this map is not injective—for example, $\pi(0)=\pi(00)=\pi(000)=0$. However, if we restrict it to words of a fixed length, then it is. In particular, the restrictions $\pi:\mathscr{U}_{n}\rightarrow\mathbb{Z}\cap[0,2^{n-1}]$ are injective maps. This gives a natural ordering on $\mathscr{U}_{n}$ and allows us to ask questions about how Ulam words are distributed. For example, we might ask about the distribution of the gaps—differences between consecutive Ulam words, interpreted as integers.
Conjecture 1.2.
$p_{n}:\mathbb{Z}_{\geq 1}$ $\rightarrow[0,\infty)$
$g$$\mapsto\frac{\#\left\{i\middle|u_{i+1}-u_{i}=g\right\}}{k_{n}-1}.$
This has a natural interpretation as a probability measure. As $n\rightarrow\infty$, the functions $p_{n}$ converge pointwise to a probability measure $p:\mathbb{Z}_{\geq 1}\rightarrow[0,\infty)$. Furthermore, let $\mu_{g}(n)$ be the mean of the probability measure $p_{n}$. Then $\mu_{g}(n)=\Theta(n^{3/10})$—indeed, it may be that there is a constant $c\approx 1.9$ such that $\mu_{g}(n)=cn^{3/10}+o(1)$.
This conjecture is well-supported by our available data—see Section 5 for details, illustrations, and further odd properties of the apparent distribution. What is interesting about this statement about average gaps is that, if true, it immediately implies Conjecture 1.1.
Theorem 1.4.
As $n\rightarrow\infty$, $\rho(n)^{-1}\asymp\mu_{g}(n)$. Consequently, Conjecture 1.2 implies Conjecture 1.1.
This is salient, since our numerical evidence for Conjecture 1.2 is arguably much stronger than for Conjecture 1.1! Again, see Section 5 for details. Finally, in Section 6, we ask the question of how $\pi(\mathscr{U}_{n})$ is distributed modulo $N$.
@@
As we discuss in Section 6, while this conjecture is consistent with the available data, it is somewhat surprising. For one thing, $\rho_{5,6}(1)=\rho_{5,6}(2)=\rho_{5,6}(3)=0$, and it takes some time before it appears to start to converge to $1/6$. For another, there is an apparent bias modulo $6$ in the distribution of the gaps.
Our code and some of our data can be found on GitHub [^1] https://github.com/asheydva/Ulam-Words.git, but it is far from efficient—as was pointed out to us Tomás Oliveira e Silva, it is possible to use bitmaps to make these computations much faster; a good implementation should give a $O(2^{n}\log(n))$ running time. However, we leave this as material for future work.
@@ 2. Definitions and Visualizations
We start with some basic definitions and constructions. Given a word $w\in S[\{0,1\}]$, we define its complement $\hat{w}$ to be the word with every instance of 00 replaced with a $1$, and vice versa. We also define the reverse $\overline{w}$, which is the word obtained by reversing the order of the letters. It was shown by Bade et al. [1] that $w\in\mathscr{U}$ if and only if $\hat{w}\in\mathscr{U}$, if and only if $\overline{w}\in\mathscr{U}$.
To better visualize the set $\mathscr{U}$, we made use of heat maps, which depict each Ulam word as a colored bar and stacks all of the words vertically—that is, for a given word, a 00 corresponds to a rectangle of one color, and a $1$ corresponds to a rectangle of a second color. An example is provided in Figure 4. In general, we abridge such diagrams: we created figures only using all the Ulam words that started with zero, since Ulam words are closed under complements. Moreover, we impose the ordering discussed in the introduction, defining $w\leq w^{\prime}$ if and only if $\pi(w)\leq\pi(w^{\prime})$. Using this way of visualizing Ulam words allows us to easily see that there is both a clear binary tree structure that governs the existence of Ulam words, as well as a chaotic element to the set where the binary tree breaks down.
@@
We can be more specific about our meaning regarding this breakdown: since Ulam words are preserved under the reverse map, this is equivalent to saying that for any $n$ there exists $n>\ell_{n}>0$ such that all possible subwords of length $\ell_{n}$ occur as the final $\ell_{n}$ characters of words in $\mathscr{U}_{n}$. In turn, that is equivalent to saying that the quotient map $\pi(\mathscr{U}_{n})\rightarrow\mathbb{Z}/2^{\ell_{n}}\mathbb{Z}$ is surjective. Our observation is that $\ell_{n}$ appears to increase as a function of $n$, albeit not very quickly—see Figure 5. Assuming that Conjecture 1.3 is true, then it would follow immediately that $\ell_{n}\rightarrow\infty$ simply by considering the case where $N=2^{\ell_{n}}$—indeed, the heat maps were the original impetus for our equidistribution conjectures. On the other hand, the “chaotic” latter half of the heat map is more of a mystery.
@@
3. Lower Bounds on Growth
Our goal in this section is to prove our lower bounds on #$\#\mathscr{U}_{n}$; we begin with Theorem 1.1, for which we need some explicit examples of Ulam words. The first three are due to Bade et al.
Theorem 3.1 (Corollary 3.4 in [1]).
There are $G(n-1)$ Ulam words of length $n$ of the form $0^{a}10^{b}$, where $G(n)$ is the $n$-th entry in Gould’s sequence.
Remark 3.1.
Gould’s sequence $G(n)$ is the number of odd entries in the $n$-th row of Pascal’s triangle; equivalently, $G(n)=2^{\#_{1}(n)}$, where #$\#_{1}(n)$ is the number of non-zero bits in the binary representation of $n$.
@@ 3. Lower Bounds on Growth
Theorem 3.2 (Theorem 3.5 in [1]).
For any $a,b\in\mathbb{Z}_{\geq 0}$, $0^{a}1^{2}0^{b}\in\mathscr{U}$ if and only if the length of the word is odd (that is, $a+b=1\pmod{2}$).
Theorem 3.3 (Theorem 3.6 in [1]).
For any $a,b\in\mathbb{Z}_{\geq 0}$ such that $a+b\geq 2$, $0^{a}1010^{b}\in\mathscr{U}$ if and only if the length of the word is even (that is, $a+b=1\mod 2$).
@@
For any $a,b\in\mathbb{Z}_{\geq 0}$ such that $a+b\geq 1$, $0^{a}1^{4}0^{b}\in\mathscr{U}$ if and only if $a+b=1\pmod{4}$.
@@
We will use proof by induction on the length of the word $n$, where the base cases $n=5,6,7,8$ can be verified directly. Assume the statement holds for all words of length strictly less than $n$, and consider the word $u=0^{k}1^{4}0^{l}$ of length $n$, where, since $n\geq 9$, at least one of $k\geq 3$ or $l\geq 3$. By applying the reverse map to switch $k$ and $l$ if necessary, we may assume that $k\geq 3$.
@@
With this, we are ready to give a proof of the general, linear bound.
Proof of Theorem 1.1.
We split into three cases.
Case 1:
$n$ is even.
By Theorem 3.3, we know that $0^{a}1010^{n-a-3}\in\mathscr{U}_{n}$ for all $30\leq a\leq n-$ 3—this yields $n-2$ Ulam words. By Theorem 3.1, we also know that there are $G(n-1)$ Ulam words of length $n$ of the form $0^{a}10^{b}$. Note that these two sets of Ulam words do not intersect (they have different numbers of ones), and since $n-1$ is odd, $G(n-1)\geq 2^{2}=4$. In total, this yields $n+2$ Ulam words.
Note that $\widehat{0^{a}1010^{n-a-3}}=1^{a}0101^{n-a-3}=0^{a_{1}}10^{b_{1}}$ if and only if $n=3$, so the reverses of the constructed Ulam words are also distinct Ulam words. Therefore, we have at least $42n+$ 4 Ulam words in this case.
Case 2:
$n=3\pmod{4}$.
By Theorem 3.2, we know that $0^{a}1^{2}0^{n-a-2}\in\mathscr{U}_{n}$ for all $20\leq a\leq n-$ 2—this yields $n-1$ Ulam words. Since $n-1=2\pmod{4}$, $G(n-1)\geq 2^{2}=4$, and so we can again use Theorem 3.1 to conclude that there are at least $4$ words $0^{a}10^{b}$ of the right length. In total, this yields $n+3$ Ulam words.
Note that $\widehat{0^{a}1^{2}0^{n-a-2}}=1^{a}0^{2}1^{n-a-2}=0^{a_{1}}10^{b_{1}}$ if and only if $a=0$ and $a_{1}=2$. Therefore, the reverses of our two families of constructed Ulam words intersect, but only in two places; therefore, we have $42(n+3)-2=2n+$ 4 Ulam words.
Case 3:
$n=1\pmod{4}$.
As in the previous case, we have $n-1$ words of the form $0^{a}1^{2}0^{n-a-2}$, but it is possible that $G(n-1)=2$, so we have to argue differently: specifically, we use Lemma 3.4 to conclude that $0^{a}1^{4}0^{n-4-a}\in\mathscr{U}$ for all $40\leq a\leq n-$ 4, which yields another $n-3$ Ulam words, for a total of at least $42n-$ 4.
Observe that $\widehat{0^{a}1^{2}0^{n-a-2}}\neq 1^{a}0^{2}1^{n-a-2}=0^{a_{1}}1^{4}0^{n-4-a_{1}}$ ever, so we may simply double our count of Ulam words. In total, we have $84n-$ 8, which is at least $42n+$ 4 if $n\geq 6$.
In each case, we have identified at least $42n+$ 4 distinct Ulam words. ∎
Next, we tackle the exponential bound, which we approach in a completely different fashion using the following lemma.
Lemma 3.5.
For any $n\in\mathbb{Z}_{\geq 1}$,
#$\#\mathscr{U}_{n}^{2}\leq\#\mathscr{U}_{n}+\#\mathscr{U}_{n+1}+\ldots+\#\mathscr{U}_{2n}.$
Proof.
Consider the set
$X:=\left\{(w_{1},w_{2})\in\mathscr{U}_{n}^{2}\middle|w_{1}\neq w_{2}\right\}.$
For any $(w_{1},w_{2})\in X$, either $w_{1}{}^{\frown}w_{2}\in\mathscr{U}_{2n}$ or there exists $v_{1}\in\mathscr{U}_{k}$, $v_{2}\in\mathscr{U}_{2n-k}$ such that $w_{1}{}^{\frown}w_{2}=v_{1}{}^{\frown}v_{2}$, where $k\in[1,n-1]\cup[n+1,2n]$; of course, if $k\in[1,n-1]$, then $2n-k\in[n+1,2n]$, and so we may conclude that
#$\#X\leq\#\mathscr{U}_{n+1}+\ldots+\#\mathscr{U}_{2n}.$
On the other hand,
#$\#X=\#\mathscr{U}_{n}^{2}-\#\mathscr{U}_{n}.$
As a consequence, we get the following very weak lower bound: for any $n\in\mathbb{Z}_{\geq 1}$,
(1) $\max_{n\leq i\leq 2n}\#\mathscr{U}_{i}\geq\frac{\#\mathscr{U}_{n}^{2}}{n+1}.$
This is sufficient for our purposes.
Proof of Theorem 1.2.
Choose any $n_{1}\in\mathbb{Z}_{\geq 6}$ and let
$\alpha_{0}:=\left(\frac{\#\mathscr{U}_{n_{1}}}{2(n_{1}+1)}\right)^{1/n_{1}}.$
Then #$\#\mathscr{U}_{n_{1}}=2(n_{1}+1)\alpha_{0}^{n_{1}}$. By Theorem 1.1, we know that #$\#\mathscr{U}_{n_{1}}>2n_{1}+2$, hence $21<\alpha_{0}<$ 2. Ergo, for any $1<\alpha<\alpha_{0}$, #$\#\mathscr{U}_{n_{1}}=2C(n_{1}+1)\alpha^{n_{1}}$ for some $C>1$. Recursively define a sequence $n_{1},n_{2},\ldots$ such that $n_{i}$ is the unique integer $n_{i-1}\leq n_{i}\leq 2n_{i-1}$ such that #$\#\mathscr{U}_{n_{i}}$ is as large as possible. We’ll prove by induction that #$\#\mathscr{U}_{n_{k}}\geq 2C^{2^{k}}(n_{k}+1)\alpha^{n_{k}}$. The base case $k=1$ is clear; for the induction step, we can apply our bound (1) to see that
$\mathscr{U}_{n_{k+1}}$ $=\max_{n_{k}\leq i\leq 2n_{k}}\#\mathscr{U}_{i}\geq\frac{\#\mathscr{U}_{n_{k}}^{2}}{n_{k}+1}$
$>\frac{4C^{2^{k+1}}(n_{k}+1)^{2}}{n_{k}+1}\alpha^{2n_{k}}$
$>2C^{2^{k+1}}(2n_{k}+1)\alpha^{2n_{k}}$
$\geq 2C^{2^{k+1}}(n_{k+1}+1)\alpha^{n_{k+1}}.$
Since $C^{2^{k}}\rightarrow\infty$, we can conclude that for all $c>0$, there exist infinitely many $n$ such that #$\#\mathscr{U}_{n}\geq c\alpha^{n}$. Therefore,
$\limsup_{n}\frac{\#\mathscr{U}_{n}}{\alpha^{n}}=\infty,$
as was claimed. All that remains is to find a value of $\alpha_{0}$ that works. In our case, we used $n_{0}=30$, which gives the $\alpha_{0}$ in the statement of the theorem. ∎
As can be seen from the proof of this theorem, if it is true that #$\#\mathscr{U}_{n}=\Theta(n^{-3/10}2^{n})$ as we conjecture, then it will be possible to improve the constant $\alpha_{0}$ arbitrarily close to $2$ simply by computing #$\#\mathscr{U}_{n}$ for larger and larger $n$. Unfortunately, this rapidly becomes impractical and in any case will never suffice to prove the theorem with $\alpha_{0}=2$, as we suspect must be true.
4. Patterns in Ulam Words
Bade et al. proved various results regarding Ulam words containing certain patterns [1]. We offer a similar collection of results. We begin with a couple of intermediate results that allow us to prove Theorem 1.3.
Theorem 4.1 (Theorem 3.4 in [1]).
A word of the form $0^{a}10^{b}$ is Ulam if and only if
$\left(\begin{smallmatrix}a+b\\ a\end{smallmatrix}\right)=1\mod 2.$
Lemma 4.2.
Proof.
  1. (1)
    $1{}^{\frown}1^{2}0^{n-3}$ and
  2. (2)
    $1^{3}0^{n-4}{}^{\frown}0$,